# Confidential Prompting: Protecting User Prompts from Cloud LLM Providers

### v2

@@ 2.3.1 Chaffing and Winnowing
In native chaffing and winnowing, a sender secures sensitive messages by mixing them with fake ones, known as chaffing. For example, to secure the message “Meet at 5pm”, the sender might transmit it along with variations like “Meet at $X$pm”, where $X$ ranges from 1 to 12. Assuming both the sender and the receiver share a symmetric secret key for signing and verifying the message authentication code (MAC) carried with message, the receiver can authenticate whether a message originates from the sender. Instead of appending an authentic MAC with a fake message, the sender includes a fake MAC with each fake message. This enables the receiver to identify and discard fake messages, known as winnowing.
The premise of confidentiality without encryption in the above protocol is that (1) an eavesdropper cannot distinguish between authentic and fake messages based on their content, while (2) the receiver can. This is achieved by (1) generating fake messages (and corresponding fake MACs) that appear authentic, and (2) using a shared symmetric secret key to differentiate the authentic MAC from the faked ones.
2.3.2 Security Model
The chaffing and winnowing concept described above achieves a certain level of confidentiality without encryption. However, naively applying such technique to messages in natural language is not information-theoretically secure because the distribution of the generated messages may not be uniform. The key challenges are how to generate fake messages that appear authentic and how to quantify their “authenticity”. For example, in the phrase “I am a teenager. I am $X$ years old.”, a poorly designed chaffing might consider $X$ to be any positive integer to maintain the grammatical correctness. However, a more effective chaffing should realize that $X$ is most likely between 13 and 19, and it would be highly improbable for $X$ to exceed 30, given the context of “teenager”.
To quantify the “authenticity” of fake messages, we introduce the concept of $\epsilon$-close over a distribution $\mathcal{D}$ to bound the “distance” between two sets of messages:
Definition 2.1 ($\epsilon$-close over $\mathcal{D}$).
In this paper, unless otherwise specified, $\mathcal{D}$ refers to the true probability distribution of messages, which is inaccessible to anyone. A language model $LM$ can approximate $\mathcal{D}$, denoted as $\mathcal{D_{\text{LM}}}$. We define the “distance” between two distributions over a set of messages $U$ as
If $\Delta_{U}(\mathcal{D},\mathcal{D_{\text{LM}}})$ is close to zero, we say the language model $LM$ provides a sufficiently accurate approximation of $\mathcal{D}$ over the subdomain $U$.
Intuitively, if there are $\lambda$ fake messages that are $\epsilon$-close to the authentic message over $\mathcal{D}$ and $\epsilon$ is small, these fake messages appear authentic, as their probabilities are close to that of the authentic message, bounded by $\epsilon$. In the case that $\epsilon$ closes to zero, the distribution across these $\lambda+1$ messages will be nearly uniform. In other words, even if an adversary obtains the plaintext of these $\lambda+1$ messages, they can only guess randomly, with a probability of approximately $\frac{1}{\lambda+1}$ of being correct, and an advantage bounded by $\epsilon$. However, since $\mathcal{D}$ is inaccessible, we approximate it with $\mathcal{D_{\text{LM}}}$ and assume that this approximation is sufficiently accurate. In §5.3, we provide a formal analysis showing that the total error is bounded by both $\epsilon$ and $\Delta_{U}(\mathcal{D},\mathcal{D_{\text{LM}}})$, where $U$ represents the union of the authentic and fake message sets.
We note that both $\epsilon$ and $\lambda$ are key security factors. Our design of prompt obfuscation allows users to configure the desired values for these security factors, that is, $\epsilon_{0}$ and $\lambda_{min}$. If fewer than $\lambda_{min}$ qualified candidates that are $\epsilon_{0}$-close (e.g., $\lambda$ is less than 8 in the example of “I am a teenager. I am $X$ years old.”), prompt obfuscation will alert the user to a possible information leak. Details can be found in §5.
@@ 2.4 Related Work
We provide a brief overview of the existing literature on privacy-preserving LLM inference, particularly in the context of protecting information in prompts. Beyond the CVM-based solutions presented in §2.2 and Figure 2, research in this area has predominantly focused on three main approaches: differential privacy (DP), multiparty computation (MPC), and homomorphic encryption (HE).
Several studies have applied the DP principles to in-context learning by injecting noise into token distributions, generating few-shot random examples, or tuning prompts [32, 6, 5, 33]. However, these DP-based methods are task-specific (i.e., applicable only for group-level information) and compromise output invariance. Although our approach shares the DP objective of preventing observers from inferring individual user information from the overall output, SMD and PO are lossless, maintaining output invariance.
MPC has been widely applied to protect input privacy in DNN inference, and its use in LLM inference has been explored, such as using MPC-based matmul or offloading some Transformer layers [9, 34]. Despite its popularity, existing MPC methods compromise model confidentiality and are vulnerable to reconstruction attacks. HE, often combined with MPC to provide end-to-end privacy guarantees, has also been applied to DNN inference [35, 10] and, more recently, to LLM [36, 37]. However, these methods remain impractical for widespread use, not only due to their significant computational overhead but also limitations in supporting nonlinear computations. As a result, nonlinear functions are often replaced with polynomial approximations, which impedes the use of existing well-trained models and reduces accuracy compared to state-of-the-art models.
A more recent approach focuses on privacy-preserving prompt engineering [7, 38], which anonymizes sensitive information in prompts by substituting sensitive segments with labels or alternative tokens before providing the entire prompt to the LLM. This approach shares similarities with our prompt obfuscation design, as both involve replacing segments in prompts. However, anonymization in this approach alters the output and reduces accuracy. In contrast, OSMD not only ensures output invariance, but also adds an additional layer of protection by retaining the entire prompt (excluding public segments such as system messages) and the private KV cache within the CVM. To the best of our knowledge, OSMD is the first scalable design that applies confidential computing to LLM serving without trusting the LLM itself.
@@ 3 Design Overview
In this section, we introduce the design goals of Confidential Prompting (§3.1), its threat model (§3.2), and the overview of OSMD (§3.3).
3.1 Design Space
In addition to the primary goal of securing users’ prompts in cloud-hosted LLM serving, we consider several constraints and trade-offs in the design of Confidential Prompting.
Constraints  Our design respects three key constraints:
@@ 3.1 Design Space
  • Compute efficiency: Protection measures must not significantly increase the cost of the service. Although security is not free, we believe that a more affordable solution will be more attractive to users.
Trade-offs  We make two important trade-offs in designing Confidential Prompting. First, although it is desirable to achieve prompt confidentiality and compute efficiency simultaneously, when faced with a choice, we prioritize confidentiality over efficiency, since protecting prompt confidentiality is our main objective. Second, we strike a balance between scalability in terms of the number of users and throughput of the prompting for an individual user. Although it is preferable to maximize both, when we have to choose one over the other, we prioritize scalability, as cloud LLMs typically serve a large number of users.
3.2 Threat Model
3.2.1 Trusted Computing Base (TCB)
@@
We assume that the users’ local machines and the cloud’s computing hardware, such as CPUs and GPUs, are trustworthy. Particularly we trust the hardware’s architectural extensions related to confidential computing, such as Intel TDX [26], AMD SEV [25] and Nvidia GPU CC [27]. This means that users can verify the integrity of their own Confidential Virtual Machines (CVMs) running in the cloud and once verified, these CVMs are trusted. Consequently, we treat a user and its CVM as the same party, even though the CVMs are hosted in the cloud.
@@
We assume that the communication channels from users to their CVMs are secure, for example, encrypted with TLS. On the other hand, because CVMs are hosted in the cloud, the cloud vendor can control the outbound network bandwidth allocated to the CVMs, and even inspect the outbound data. This bandwidth management and inspection are crucial to maintaining model confidentiality while allowing the CVMs to read the LLM weights.
@@
Finally, we assume the users and their CVMs privately share the key (or called seed) for a pseudorandom function.
3.2.2 Threat Model
We adopt the Honest-but-Curious (HBC) adversary model with respect to cloud LLM providers. In this model, the cloud LLM provider faithfully follows the prescribed computation steps for LLM inference, but seeks to extract as much information as possible from users. Specifically, we assume that the cloud LLM provider has access to the tokens generated during the decoding phase and their associated public KV attention states. In contrast, user prompts and the associated private KV attention states remain securely within the CVMs, inaccessible to the cloud LLM provider (2(d)). However, the cloud LLM provider may attempt reconstruction attacks to retrieve sensitive information stored in the CVMs, leveraging the known public states and its observation of the interactions between the LLM and the CVMs.
@@ 3.2.2 Threat Model
We do not consider attacks that can compromise the CVMs, the communication channels, or the pseudorandom function (PRF). Denial of service (DoS) attacks are also out of scope in this work. We note that confidential computing systems generally assume stronger threat models than HBC. Confidential Prompting requires a weaker threat model as it offloads most computations outside of the CVM for efficiency, but does not check the integrity of their results. We discuss possible strategies for a stronger threat model in §8.
3.3 OSMD Overview
The objective of Confidential Prompting is to protect the user’s input, or prompts, when interacting with the cloud LLM. OSMD’s core design components, Secure Multi-party Decoding (SMD) and Prompt Obfuscation (PO), are designed to satisfy all the constraints outlined in §3.1. Figure 1 illustrates how SMD and PO work together to defend against the threat model described in §3.2. We next provide a detailed explanation of the collaboration between SMD and PO.
Initialization  The user first initializes a CVM in the cloud and verifies its integrity through remote attestation (§2.2). Next, the user tags the sensitive portions of their prompt, encrypts it locally, and transmits the encrypted prompt to the CVM. Assuming that the user and the CVM share symmetric secret keys for a secure channel, the CVM can exclusively decrypt and parse the encrypted prompt. In particular, tagging sensitive portions can be automated using tools such as Google’s PII detection tool, as demonstrated in our prototype §7.1 and discussed in §8.
@@ 3.3 OSMD Overview
Prefill and Prompt Obfuscation (§4.1 and §5)  For each tagged sensitive subsequence $S_{target}$ in the prompt $S$, the CVM samples a set of fake n-grams $S_{fake}$5.1), such that for each $C\in S_{fake}$,
$\lVert P(C|S-S_{target})-P(S_{target}|S-S_{target})\rVert<\epsilon$
where $\epsilon$ is an error bound parameter close to 0, and the operation $S_{a}-S_{b}$ returns the result of removing subsequence $S_{b}$ from $S_{a}$. The CVM generates $\lambda$ virtual prompts using $S_{fake}$ and generates idx with a pseudorandom function (PRF), which serves as the index of the authentic prompt among the $\lambda+1$ prompts. Finally, the CVM computes the private KV cache for all $\lambda+1$ prompts.
As its name suggests, virtual prompts are “virtual”. Instead of computing the KV cache for each of the $\lambda$ virtual prompts individually, the CVM uses a memory efficient approach to joinly store them (§5.2). Additionally, allowing the CVM to access the LLM’s weights during the prefill phase does not compromise the model confidentiality constraint, as discussed in §3.2 and §4.1.
Secure Multi-party Decoding (§4.2)  SMD decouples the attention computation in the decoding phase into a two-party computation. One party, the CVM, has exclusive access to the user prompts and maintains the private KV cache within the CVM. SMD enables scalable use of CVMs, as the CVMs do not require access to LLM weights after the prefill phase, addressing the compute efficiency challenges of applying CVMs to the LLM service. Furthermore, SMD is lossless, ensuring that it meets the output invariance constraint.
@@
Although SMD is vulnerable to reconstruction attacks, PO mitigates this issue at the cost of redundant decoding and communication. After generating output tokens, the CVM sends the responses for all $\lambda+1$ prompts back to the user. It is worth noting that the user already knows the value of the index idx, because we assume the user and the CVM share the key (or called seed) of the PRF (§3.2). This allows the user to perform winnowing and extract the authentic response.
Offloading Sampling to User  As a variant of the Prompt Obfuscation described above, users may choose to offload the fake n-gram sampling to their local devices. Our sampling algorithm (Algorithm 1) requires a language model $LM$ to approximate the authentic language distribution $\mathcal{D}$. Sampling in the CVM can leverage either the cloud LLM or a small language model (SLM) hosted within the CVM, although the latter may increase the CVM size. In contrast, offloading is practical and beneficial because: (1) According to our analysis in §5.3, the security of PO depends on the distance $\Delta_{U}(\mathcal{D},\mathcal{D}_{\text{LM}})$, which is independent of whether the model used is a powerful cloud LLM or a less powerful SLM on the user’s device. In fact, a local SLM fine-tuned with user data may be more effective, as it can automatically detect and better anonymize sensitive information based on the user’s habits, and thus more secure than generic cloud LLMs. In practice, local models such as Apple Intelligence are gaining popularity. (2) The tagged sensitive subsequences $S_{target}$ are usually “common” subtexts, which do not necessarily require a powerful cloud LLM for sampling replacements. (3) Offloading reduces the size of the CVM and the time to sample fake n-grams on the server side.
Further discussion of the variants of PO is provided in §8. In the following, we introduce the design details without offloading, as the modifications for the offloading scenario are straightforward.
4 Efficient Prompt Protection with Confidential Computing
@@
The main obstacles to using CVMs for LLM serving stem from keeping LLM weights in each CVM, as discussed in §2.2. We solve this challenge by eliminating the need to store LLM weights in CVMs after the initial prefill operation. Our key insight is that the decoding operation in LLMs can be formulated as a secure multi-party computation, where only one party needs to access the LLM weights. Specifically, we distinguish $K,V$ attention states into private and public parts. The user prompt’s $K,V$ states are private and are kept confidential in the CVM, while the generated sequence’s $K,V$ states are public and shared with the LLM. We detail the inference protocol as follows.
We define three participants: (1) Alice, the user; (2) Eve, the cloud LLM provider; and (3) Ted, Alice’s CVM instance running on Eve’s premises. As mentioned in §3.2, we treat Alice and Ted as one party, assuming their communication is secure, while Eve follows the protocol, but tries to extract the maximum information from Ted.
@@
    Prefill4.1): Alice sends a prompt to Ted. Ted computes ($K_{\mathrm{Pvt}},V_{\mathrm{Pvt}}$) states for the prompt and generates the first token. Ted sends the first token to Eve.
@@
    Decoding4.2): For each transformer layer, Eve computes $Q_{\mathrm{New}},K_{\mathrm{New}},V_{\mathrm{New}}$ of the received token, appends $K_{\mathrm{New}},V_{\mathrm{New}}$ to $K_{\mathrm{Pub}},V_{\mathrm{Pub}}$, and shares $Q_{\mathrm{New}}$ with Ted. Ted responds with $\sigma(Q_{\mathrm{New}}K_{\mathrm{Pvt}}^{\top})V_{\mathrm{Pvt}}$. Eve computes $\sigma(Q_{\mathrm{New}}K_{\mathrm{Pub}}^{\top})V_{\mathrm{Pub}}$ and recovers $\sigma(Q_{\mathrm{New}}K^{\top})V$ using Theorem 4.1. If it is the final layer, the output $Y$ is sent to Ted.
@@
    Ted samples a new token from $Y$ and sends it to Eve and Alice. Go to Step 2 until [EOS].
This protocol can be generalized to scenarios with multiple users, each with their own CVM operating in the cloud.

### v3

@@ 2.3 Chaffing and Winnowing
In an information-theoretic sense, a message is secure if an eavesdropper can do no better than guessing randomly to determine its content [50, 33]. Modern encryption achieves this by transforming messages into ciphertexts that are indistinguishable from random noise; in doing so, unfortunately, it renders the text meaningless to an LLM. In contrast, chaffing and winnowing [48, 5] secures message confidentiality without encrypting the message itself. In naïve chaffing and winnowing, a sender secures sensitive messages by mixing them with fake ones, known as chaffing. For example, to secure the message “Meet at 5pm”, the sender might transmit it along with variations like “Meet at $X$pm”, where $X$ ranges from 1 to 12. Assuming both the sender and the receiver share a symmetric secret key for signing and verifying a message authentication code (MAC) sent along with a message, the receiver can authenticate whether a message originates from the sender. The sender includes a fake MAC with each fake message, which enables the receiver to identify and discard fake messages, known as winnowing.
@@ 2.4 Related Work
In recent years, researchers have explored various approaches beyond confidential computing to preserve the privacy of user prompts in LLM inference [14].
Differential Privacy protects prompt confidentiality by injecting noise into token distributions [59, 42], generating few-shot random examples [54], or tuning prompts [21]. However, these methods are task-specific and compromise output invariance. In contrast, SPD and PO maintain output invariance.
Multi-Party Computation (MPC)-based methods utilize secret sharing that cryptographically splits a number $x$, either an LLM weight or a prompt token, into multiple numbers, i.e., $x=x_{1}\oplus x_{2}\oplus...\oplus x_{n}$. Then they distribute each split $x_{i}$ to an untrusted party. The user then derives the LLM inference output by crytographically combing the outputs of these parties. This technique suffers from multiple problems. First, the compute parties must not collude. Second, secret sharing is not efficient for all LLM operations. Recognizing its inefficiency, the authors chose to modify the model, e.g., using ReLU instead of SoftMax [2], or use a much smaller model distilled specially [27], violating output invariance.
Homomorphic encryption (HE) enables computations on encrypted data and is often combined with MPC to guarantee end-to-end privacy in LLM inference [31, 43, 19, 9]. However, its significant overhead impedes its use in real-world applications, particularly for supporting nonlinear functions. Recent works [31, 43, 9] replace these functions with approximations, further impeding the use of existing well-trained models.
To our knowledge, other approaches [29, 13] do not protect the prompts from untrusted cloud hosted LLMs. For example, the ensemble-only method [13] mitigates the membership inference attacks (MIAs) assuming the attackers have access only to the LLM’s prediction outputs. By the time of writing, OSPD is the first scalable design that applies confidential computing to LLM serving without trusting the LLM.
@@ 3 Design Overview
We introduce the design goals (§3.1) and threat model (§3.2) of Confidential Prompting, and an overview of OSPD (§3.3).
3.1 Design Goals
Our primary goal is to secure user prompt confidentiality in cloud-hosted LLM serving. Depending on whether the attackers have unlimited computational resources or not, we aim to protect the sensitive segments in user prompts, such as Personally Identifiable Information (PII), or the entire prompts. Beyond user prompt confidentiality, we target at another three goals in designing Confidential Prompting.
@@ 3.1 Design Goals
  • Compute efficiency: Protection measures must not significantly increase the cost. Though security is not free, we believe a more efficient approach is more attractive to users.
3.2 Trust and Threat
@@
Trusted Computing Base (TCB)  We assume the users’ local machines and the cloud’s computing hardware, such as CPUs and GPUs, are trustworthy. Particularly we trust the confidential computing related hardware’s architectural extensions, such as Intel TDX [23], AMD SEV [3] and Nvidia GPU CC [40]. This means that users can verify the integrity of their own Confidential Virtual Machines (CVMs) running in the cloud and once verified, these CVMs are trusted. Consequently, we treat a user and its CVM as the same party, even though the CVMs are hosted in the cloud.
@@
We assume the communication channels between users and their CVMs are secure, for example, encrypted with TLS. However, as CVMs are hosted in the cloud, the cloud vendor can control the outbound network bandwidth allocated to CVMs, and inspect the outbound data. This bandwidth management and inspection are crucial to maintain model confidentiality while allowing the CVMs to read LLM weights.
@@
Finally, we assume the users and their CVMs privately share the key (or seed) for a pseudorandom function (PRF).
Threat Model  We consider the cloud vendor and the LLM provider as the same party, so called cloud LLM providers. We assume the users and the cloud LLM providers are mutually untrusted and seek to uncover other’s secrets. We adopt two variants of Honest-but-Curious (HBC) adversary model with respect to the cloud LLM providers. The only difference between the two variants is if the cloud LLM providers have (and are willing to use) unlimited computational resources to perform an attack. In these HBC variants, the cloud LLM providers faithfully follow the prescribed computation steps for LLM inference, but seek to extract user secrets in prompts.
We do not consider attacks that compromise the CVMs, the communication channels, or the pseudorandom function (PRF). Consequently, user prompts and the private KV cache remain securely within the CVMs, inaccessible to the cloud LLM provider (2(d)). However, under the HBC variant that the cloud LLM providers have unlimited computational resources, the providers may attempt reconstruction attacks to retrieve sensitive information stored in the CVMs.
@@
Denial of service (DoS) attacks are out of scope in this work. We note that confidential computing systems generally assume stronger threat models than HBC. Confidential Prompting assumes a weaker threat model as it offloads most computations out of the CVMs for efficiency. We discuss possible strategies for a stronger threat model in §A.4.
3.3 OSPD Overview
Confidential Prompting aims to achieve all goals outlined in §3.1. As illustrated in Figure 1, its core design components, Secure Partitioned Decoding (SPD) and Prompt Obfuscation (PO), collaborate to defend against the threat model in §3.2. We next provide a detailed overview of this collaboration.
Initialization  To use OSPD for confidential prompting, a user creates its CVM on cloud in the same way as how a CVM is initialized in common cases. That is, the user first creates its CVM with a prototype of OSPD (§6) in the cloud, and verifies the CVM integrity through remote attestation. Next, the user encrypts its prompt locally before transmitting the prompt to the CVM, which can decrypt the prompt exclusively. During the entire setup, all communications are secure (§3.2).
Prompt Obfuscation5)  PO has two steps to generate virtual prompts: (i) tagging sensitive segments in a prompt and (ii) sampling fake n-grams for each tagged segment.
For (i), a user can deploy any existing tools in its CVM to identify and tag sensitive segments in its prompt. The user can also train a small language model (SLM) for the tagging. For simplicity, we use Google’s PII detection tool to tag Personally Identifiable Information (PII) in our prototype (§6) as an example implementation.
@@ 3.3 OSPD Overview
For (ii), the CVM processes each tagged sensitive segment $S_{target}$ in the prompt $S$ independently (§5.1). That is, the CVM samples a set of fake n-grams $F$, such that for each $S_{fake}\in F$,
$\lVert P(S_{fake}|S-S_{target})-P(S_{target}|S-S_{target})\rVert<\epsilon$
where $\epsilon$ is an error bound parameter close to 0, and the operation $S_{a}-S_{b}$ returns the result of removing segment $S_{b}$ from $S_{a}$. As for the probability function $P$, it returns the probability of a token sequence, modeled by a language model (LM).
The CVM generates $\lambda$ virtual prompts with $F$ and a random number idx with a pseudorandom function (PRF), serving as the index of the authentic prompt among $\lambda+1$ prompts.
Prefill4.1)  The CVM computes private KV cache for all $\lambda+1$ prompts, always keeping these $\lambda+1$ prompts and their private KV cache within the CVM without any exposure.
We note that the CVM requires LLM weights for the computation during prefill. Once the CVM gets the LLM weights from the cloud LLM providers, it is not allowed to send any data out of the cloud, except the output tokens in plaintext generated by the cloud LLM instance. The cloud vendor limits its network bandwidth and inspect all outbound data to avoid LLM weights leakage, thereby ensuring model confidentiality.
Secure Partitioned Decoding4.2)  SPD partitions the attention score computation during decode phase into a two-party computation. That is, the CVM and the cloud LLM computes the private ($A_{pvt}$) and the public ($A_{pub}$) attention score respectively. Specifically, the private attention score $A_{pvt}$ is computed with the private KV cache, which is already precomputed during prefill and stored in the CVM. Consequently, the CVM does not require the LLM weights for computation during decode, thereby remaining a small footprint in decode and improving the scalability.
After computation, the CVM sends $A_{pvt}$ to the cloud LLM, which then merges the $A_{pvt}$ with the $A_{pub}$ to get the full attention score. As the attention score decomposition is lossless, SPD allows the cloud LLM to generate tokens without compromising output invariance.
Additionally, SPD maintains compute efficiency as it allows the cloud LLM to batch process for multiple users, and the CVM memory footprint remains small during decode.
Complementarity between SPD and PO  SPD and PO are complementary techniques that collaboratively defend against attacks which would likely succeed if either technique were employed independently. Firstly, SPD alone is vulnerable to reconstruction attacks. Conversely, like many existing obfuscation methods, an attacker can improve their chances of defeating PO by obtaining detailed knowledge of the distribution of virtual prompt generation through statistical analysis.
However, these attacks can breach OSPD only at a high cost and with a low probability of success. Within the OSPD framework, all $\lambda+1$ prompts and their private KV cache securely reside within the CVM. The protection provided by SPD significantly increases the cost of collecting statistical information, such as performing reconstruction attacks to extract all $\lambda+1$ prompts and deduce distribution data. With limited statistical information under a certain amount of computational resources, an attacker gains little advantage in breaching PO beyond random guessing. Consequently, even in the worst-case scenario where an adversary reconstructs all $\lambda+1$ prompts, the probability of accurately identifying the authentic prompt increases by merely $\epsilon$ compared to random guessing among the $\lambda+1$ possible prompts.
@@
Extract Authentic Response  The CVM sends the responses of all $\lambda+1$ prompts to the user. It is worth noting that the user already knows the index idx, as we assume the user and the CVM share the key of the PRF (§3.2). This allows the user to perform winnowing and extract the authentic response.
4 Efficient Prompt Protection with
Confidential Computing
@@ 4 Efficient Prompt Protection with
Keeping a dedicated copy of LLM weights in multiple CVMs leads to limited scalability as discussed in §2.2.2. This is the main obstacle to use per-user CVMs for LLM serving while untrusting the LLM provider. We solve this challenge by eliminating the need to store LLM weights in CVMs after the initial prefill phase. Our key insight is the decoding in LLMs can be formulated as a secure multi-party computation, where only one party needs access to the LLM weights. Specifically, we distinguish $K,V$ attention states into private and public parts. The user prompt’s $K,V$ states are private and are kept confidential in the CVM, while the generated tokens’ $K,V$ states are public and shared with the LLM. For simplicity, we detail the protocol with one-user scenario as follows.
There are three participants in the protocol: (i) the user; (ii) the user’s CVM; and (iii) the cloud LLM.
@@
    Prefill4.1): The user sends a prompt to its CVM, which computes $K_{\mathrm{Pvt}},V_{\mathrm{Pvt}}$ states of the prompt and generates the first token. Then the CVM sends the token to the LLM.
@@
    Decode4.2): As in Figure 3, for each transformer layer:
    1. (a)
      The cloud LLM computes $Q_{\mathrm{New}},K_{\mathrm{New}},V_{\mathrm{New}}$ of the received token, appends $K_{\mathrm{New}},V_{\mathrm{New}}$ to $K_{\mathrm{Pub}},V_{\mathrm{Pub}}$ to maintain public KV cache, and sends $Q_{\mathrm{New}}$ to the CVM.
    2. (b)
      The CVM responds with private attention score
      $A_{pvt}=\sigma(Q_{\mathrm{New}}K_{\mathrm{Pvt}}^{\top})V_{\mathrm{Pvt}}$.
    3. (c)
      The cloud LLM computes public attention score $A_{pub}=\sigma(Q_{\mathrm{New}}K_{\mathrm{Pub}}^{\top})V_{\mathrm{Pub}}$, and merge $A_{pvt}$ with $A_{pub}$ to recover full attention score $Y=\sigma(Q_{\mathrm{New}}K^{\top})V$ using Theorem 4.1.
    4. (d)
      If it is the final layer, the output $Y$ is sent to the CVM.
@@
    The CVM samples a new token from $Y$ and sends it to the cloud LLM and the user. Then go to Step 2 until [EOS].
This protocol can be generalized to scenarios with multiple users, each with its own CVM operating in the cloud.